# Resolution-Adaptive Micro-Doppler Spectrogram for Human Activity Recognition

### v2

@@ IV Performance Analysis
Figure 5: Confusion matrices for models trained on conventional and RA spectrogram datasets.
@@
Fig. 5 presents the confusion matrices that compare actual activities with those predicted by the HAR model. The ResNet18 and ConvNet-Shallow models were used with datasets constructed using conventional and proposed RA spectrograms. The training datasets consisted of P2–P5 data, whereas the test dataset consisted of P1 data. The comparative results show that the models trained on RA spectrograms achieve a higher average recognition accuracy than those trained on conventional spectrograms. In particular, there is a marked improvement in the recognition accuracy of the A1 activity (falling), which is a critical factor in HAR because of its significance in safety and health monitoring applications.
@@
Tables I and II present the average recognition accuracies for the models based on various time–frequency analysis methods. In most cases, the models that use the proposed RA spectrogram demonstrate the highest recognition accuracy. It should be noted that for the P1 test dataset, the VGG19 model achieves a recognition accuracy of 93.7% when the RA spectrogram dataset is used. This represents an improvement of 4.31%–5.64% over the same model with other time–frequency representation methods. Similarly, for the P2 test dataset, the ConvNet-Shallow model achieves an accuracy of 86.33% with the RA spectrogram, which exceeds the accuracy of other preprocessing methods by 4.78%–11.04%. These results demonstrate that the RA spectrogram provides a refined resolution that is suitable for human activity analysis. The proposed method more effectively captures and classifies micro-Doppler signature patterns compared with existing methods.
TABLE I: Accuracy of HAR models for datasets using P1 test data
Model Recognition Accuracy
Spectrogram SPWVD RIDHK RA Spectrogram
CNN[16] 88.39 86.40 86.40 91.87
ConvNet-Deep[17] 90.88 89.72 91.38 94.69
ConvNet-Shallow[17] 82.59 77.78 77.61 85.74
ResNet18[18] 87.56 92.37 88.89 94.03
VGG16[19] 89.88 89.55 91.38 93.86
VGG19[19] 89.39 88.72 88.06 93.70
TABLE II: Accuracy of HAR models for datasets using P2 test data
Model Recognition Accuracy
Spectrogram SPWVD RIDHK RA Spectrogram
CNN[16] 95.39 93.57 93.41 96.87
ConvNet-Deep[17] 96.54 95.39 94.40 97.69
ConvNet-Shallow[17] 77.92 75.29 81.55 86.33
ResNet18[18] 90.28 92.75 94.23 97.03
VGG16[19] 93.74 93.41 93.25 95.39
VGG19[19] 92.09 90.94 94.73 94.56
V Conclusion
We propose a time–frequency representation method that adaptively adjusts the resolution. The proposed method explores the frequency range of micro-Doppler signatures within a spectrogram and applies a nonlinear transformation to the frequency resolution within the identified range. Experimental results demonstrate that the proposed approach achieves a higher recognition accuracy compared with various time–frequency analysis techniques. Although this study focuses on environments in which a single activity is performed during the measurement time, continuous HAR systems may encounter idle times with no activity in the time domain. Therefore, future work will focus on enhancing the HAR system by developing methods that simultaneously adapt the resolution in the time and frequency domains.

### v3

@@ 4 Performance Analysis
Figure 5: Confusion matrices for model trained on conventional and RA spectrogram datasets.
@@
Fig. 5 presents the confusion matrices that compare actual activities with those predicted by the HAR model. The ResNet18 model was trained on datasets generated using conventional and proposed RA spectrograms. Training datasets included data from participants P2–P5, while the test dataset comprised data from participant P1. Results indicate that the model trained on RA spectrograms achieves higher average recognition accuracy than the one trained on conventional spectrograms. In particular, recognition accuracy significantly improved for activity A1 (falling), a critical factor in HAR due to its importance in safety and health monitoring applications.
Table 1: Accuracy of HAR models for various datasets
@@
Table 1 presents the average recognition accuracies for the models based on various time–frequency analysis methods. The proposed RA spectrogram consistently outperforms the conventional spectrogram, SPWVD, and RIDHK representations across all datasets (P1–P5) and under the stratified 5-fold cross-validation (CV). Based on the LOOCV results over the P1–P5 datasets, the proposed method improves the recognition accuracy by 0.9 %, 2.5 %, 3.3 %, 5.2 %, 1.7 %, and 2.6 % for CNN, ConvNet-Deep, ConvNet-Shallow, ResNet18, VGG16, and VGG19, respectively, compared with conventional spectrograms. Under the stratified CV dataset, the proposed approach yields an average improvement of 3.1 %, 3.3 %, and 2.9 % over the conventional spectrogram, SPWVD, and RIDHK representations, respectively, when averaged across all tested models. These results demonstrate that the proposed method provides a refined resolution suitable for human activity analysis, enabling more effective capturing and classification of micro-Doppler signature patterns compared with existing methods.
5 Conclusion
We proposed an adaptive time–frequency representation method that dynamically adjusts its resolution based on activity-induced variations in micro-Doppler signatures. The proposed method identifies the relevant micro-Doppler frequency regions and selectively enhances their resolution while reducing resolution elsewhere, maintaining a fixed spectrogram size to effectively represent signature characteristics. Experimental results demonstrate that the proposed approach achieves a higher accuracy compared with various time–frequency analysis techniques. Future research will focus on extending the proposed resolution adjustment method beyond the time–frequency domain to other signal representation domains.
6 Funding
This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2025-00557790).
7 Availability of data and materials
Source code and sample datasets are available in the GitHub repository:
github.com/Signal-Park/Activity-Dependent-Spectrogram